OpenStudy (anonymous):

how do you integrate (x^9)(cos(x^5))?

OpenStudy (anonymous):

\[intx ^{9}cosx ^{5}dx. \]

OpenStudy (anonymous):

Possibly integration by parts

OpenStudy (anonymous):

i realize that, but should i don't know what values to assign for u and dv. should i substitute before i even use integration by parts?

OpenStudy (anonymous):

Try x^9 as u

OpenStudy (anonymous):

((x^10)/10)(-sinx(x^5))

OpenStudy (anonymous):

What is that Nick, like the teacher says, show your work.

OpenStudy (anonymous):

okay Integrate (x^9)we get (x^9+1)/(9+1) intregrate cos(x^5) we get -sin(x^5)

OpenStudy (anonymous):

But there is a rule, the two x quantities are multiplied, so you can't integrate straight out, right?

OpenStudy (anonymous):

yeah, so you integrate by parts.

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

substituting u for x^9 doesn't seem to work chaguanas ...

OpenStudy (anonymous):

Check this I haven't done integration by parts in a while\[x ^{9}\cos x ^{5}-45x ^{5}\sin x ^{5}\]

OpenStudy (anonymous):

+C

OpenStudy (watchmath):

wolframalpha will give you the steps as well.

OpenStudy (anonymous):

hold on are you sure that the integral of cos x^5 is is -sin x^5 - try differentiating -sin x^5.

OpenStudy (anonymous):

Yeah, mine is wrong, see the wolfram watchmath put up. That was the u sub that ecollison was talking about, I missed that.

OpenStudy (anonymous):

watchmath thanks! the link was right

OpenStudy (anonymous):

But please explain the u sub, they didn't account for the x^9 fully.

OpenStudy (anonymous):

u=x^5 not x^9

OpenStudy (anonymous):

lols